OpenStudy (anonymous):

Solve x2 − 7x + 12 = 0. x = −3, x = −4 x = 3, x = 4 x = 2, x = 6 x = −2, x = −6

OpenStudy (anonymous):

B

OpenStudy (anonymous):

do you want me to explain?

OpenStudy (anonymous):

sure why not

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Set the factor '(4 + -1x)' equal to zero and attempt to solve: Simplifying 4 + -1x = 0 Solving 4 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '-4' to each side of the equation. 4 + -4 + -1x = 0 + -4 Combine like terms: 4 + -4 = 0 0 + -1x = 0 + -4 -1x = 0 + -4 Combine like terms: 0 + -4 = -4 -1x = -4 Divide each side by '-1'. x = 4 Simplifying x = 4

OpenStudy (anonymous):

Simplifying 3 + -1x = 0 Solving 3 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '-3' to each side of the equation. 3 + -3 + -1x = 0 + -3 Combine like terms: 3 + -3 = 0 0 + -1x = 0 + -3 -1x = 0 + -3 Combine like terms: 0 + -3 = -3 -1x = -3 Divide each side by '-1'. x = 3 Simplifying x = 3

OpenStudy (anonymous):

did this help?

OpenStudy (anonymous):

Yeah it did sorry my OS crashed

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

it is OK you are so welcome :)